Company ABC is a manufacturing company that provided the following information:

1. Assume the following hours of cleaning work and the total cleaning cost (mixed cost) for the six months:

Number of hours

Total cleaning cost

January

3

$1,000

February

8

$2,400

March

5

$2,000

April

6

$2,100

May

8

$2,400

June

9

$2,500

Using the high-low method, estimate the cost formula for this cleaning cost (2 marks) and discuss the drawbacks of the high-low method (2 marks).

2. In the production of July, direct materials used totaled $380. Direct labor was $370 and the manufacturing overhead was $100. The beginning work in process was $1,000 and ending work in process was $500. The beginning finished goods were $2,000 and ending finished goods was $1,000. What were total manufacturing costs incurred for the month? (2 marks) What was the cost of goods manufactured during the month? (2 marks) What was the cost of goods sold for the month? (2 marks)
Answer #1

1. High Low method:

(Highest cleaning cost - lowest cleaning cost)/(Highest hours - cleaning hours)

Ie., (2500-1000)/(9-3) = 250.

So Fixed costs,

Under highest => (2500 - 250*9) = 250

Under lowest => (1000 - 250*3) = 250

So cost formula = 250 + 250*Units.

Drawback of High Low method is:

  • The facts and figures are based on estimates, and are not accurate.
  • The tools provided by the method are limited scope unlike regression analysis.
  • The variable and fixed costs are assumed to be constant without any variations over a long period of time that may not be practically true.
Amount ($)
Direct Material 380
Direct Labor 370
Manufacturing overhead 100
A.Total manufcaturing cost 850
Add: Begining Inventory 1,000
Less: Ending Inventory (500)
Cost of Goods Manufacturing 1350
Add: Begining Finished Goods 2000
Less: Ending Inventory of finished goods 1000
Cost of goods sold 2350